Blake has a total of $12,000 to invest into two accounts. One account earns 2% simple interest and the other earns 5% simple interest. How much should be invested in each account to earn exactly $390 at the end of one year
Answer by Boreal(15235)

x at .02
12000-x at .05
.02x+.05(12000-x)=390. With simple interest it is principal *rate*1 for a year
.02x+600-.05x=390
-.03x=-210
x=$7000 at 2% ($140)
12000-x=$5000 at 5% or $250.